OpenStudy (anonymous):

A triangular playground has an angles with measure in the ratio 8:3:9. What is the measure of the smallest angle

OpenStudy (whpalmer4):

If you have angles in the ratio 8:3:9, you can find out each angle by dividing its portion of the ratio by the sum and multiplying by the sum of the interior angles of a triangle, which of course is 180 degrees. So, for the largest angle, the angle is 9 parts of (8+3+9) = 9/20, and 9/20*180 = 81 degrees. You should be able to find the other two angles and verify that they all add up to 180.